OpenStudy (dsena):

2x/x-1=?/x(x-1)(x+2) I am being asked to make an equivalent fraction with the given denominator. Any ideas how I'd do this?

OpenStudy (dsena):

\[\frac{ 2x }{ x-1 }=\frac{ ? }{ x(x-1)(x+2) }\]

OpenStudy (danjs):

recall multiplying fractions \[\large \frac{ a }{ b }*\frac{ c }{ d }=\frac{ a*c }{ b*d }\] and you can multiply a fraction by 1 and it will not change, 1 can be a ratio of anything over itself.

OpenStudy (danjs):

you can see the denominator has been multiplied by x*(x+2) , to keep the fraction the same, you must do that to the numerator also. \[\frac{ 2x }{ x-1 }*\frac{ x(x+2) }{ x(x+2) }=\]